# Response Matrix Estimation in Unfolding Differential Cross Sections

### v1

@@ 6 Application to Drell-Yan + jets events
Figures 17 and 18 show the Tikhonov-regularized solution with $\delta=10^{-7}$ and D’Agostini solution with 20 iterations, respectively. The regularization parameters ($\delta$ and $n_{iter}$) are selected empirically to best balance the bias-variance tradeoff. Since we do not have access to the ground truth, the proxy truth histogram in this case is the empirical particle-level histogram, which itself has a statistical uncertainty. However, the sample size is chosen to be large to ensure that this uncertainty remains small. The results show some similar characteristics to those discussed in Section 5. In particular, when using the D’Agostini solution with a moderate number of iterations, the differences between the different response matrix estimators are generally less pronounced than in the Tikhonov case. In contrast, the Tikhonov-regularized solution shows greater variance, especially in the high $p_{\perp}$ region.
@@
Figure 17: Tikhonov-regularized solution with $\delta=10^{-7}$.
Figure 18: D’Agostini solution with 20 iterations.
@@ 7 Discussion and conclusion
Author contributions:
Huanbiao Zhu: Writing – original draft, Writing – review and editing, Methodology, Conceptualization, Investigation, Visualization, Software. Andrea Marini: Writing – review and editing, Data curation. Mikael Kuusela: Writing – review and editing, Methodology, Conceptualization, Investigation. Larry Wasserman: Writing – review and editing, Methodology, Conceptualization, Investigation.
Funding sources
@@
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
The data, along with the code required to reproduce the results, are publicly available at: https://github.com/richardzhs/response_matrix_estimation.

### v2

@@ 6 Application to Drell-Yan + jets events
Figures 11(a) and 11(b) show the Tikhonov-regularized solution with $\delta=10^{-7}$ and D’Agostini solution with 20 iterations, respectively. The regularization parameters ($\delta$ and niterniter) are selected empirically to best balance the bias-variance tradeoff. Since we do not have access to the ground truth, the proxy truth histogram in this case is the empirical particle-level histogram, which itself has a statistical uncertainty. However, the sample size is chosen to be large to ensure that this uncertainty remains small. The results show some similar characteristics to those discussed in Section 5. In particular, when using the D’Agostini solution with a moderate number of iterations, the differences between the different response matrix estimators are generally less pronounced than in the Tikhonov case. In contrast, the Tikhonov-regularized solution shows greater variance, particularly in the high $p_{\perp}$ region, as evidenced by the larger fluctuations in the ratio plot shown in Figure 11.
@@
(a) Tikhonov-regularized solution with $\delta=10^{-7}$.
(b) D’Agostini solution with 20 iterations.
Figure 11: Comparison of unfolding methods on Drell-Yan data: (a) Tikhonov regularization and (b) D’Agostini iteration.
@@ 7 Discussion and conclusion
Acknowledgments